I am doing a free-falling experiment with five trials with a meter stick and need to calculate reaction time (s). I need to convert to both SI units and English units. My stats are

Fall Distance 1: 31cm

Fall Distance 2: 20cm

Fall Distance 3: 22cm

Fall Distance 4: 15cm

Fall Distance 5: 14cm

1. I need to calculate the reaction time for each and then the average reaction time

2.  I need to show conversion of fall distance from English to SI (or vice versa) for trial 1

3. Calculation of reaction tie for trial 1

Please show work - I am interested in learning so I can apply it to a different problem and understand. I very . much appreciate it.
Answer #1

1.

For free falling body reaction time is given by

d=(1/2)gt2

time,t=(2d/g)1/2

Where g=981 cm/s2 is acceleration due to gravity and d is falling distance.

For trail 1

d=0.31 cm

t1=[(231)/981]1/2=0.251sec

For trail 2

d=20 cm

t2=[(220)/981]1/2=0.202 sec

For trail 3

d=22 cm

t3=[(222)/981]1/2=0.212 sec

For trail 4

d=15 cm

t4=[(215)/981]1/2=0.175 sec

For trail 5

d=14 cm

t5=[(214)/981]1/2=0.169 sec

Average reaction time,tav=(t1+t2+t3+t4+t5)/5=(0.251+0.202+0.212+0.175+0.169)/5=0.202 sec

2.

fall distance for trail 1=31cm

In english system falling distance in feet

1 cm=0.0328084 ft

31 cm=310.0328084=1.0171 ft

In sI system length is in m

1ft=0.3048m

1.0171 ft=0.31m

3

Reaction time for Trail 1=0.251 sec (as calculated in part 1)